HW 2, Probability (353 s19.pdf Adobe Acrobat Reader DC File Ecit View Window Help Home Tools notes 02 crystal str... notes 03 atom posi.. HW 2, Probability x Sign In 140% [t] Share Over 50% of all engineers end up in management. An engineering firm needs project managers that understand the technical aspects of the product. A large firm is looking toward future needs A look at personnel records indicates 30% of their current engineers (who are not yet in management positions) have an electrical engineering degree, 42% have a mechanical engineering degree, and 21% have a software engineering degree, 27% of all the engineers have some management iraining. Specifically, I 1% have an electrical engineering degree and management training, 9% have a mechanical engineering degree and management training, and 2% have a software engineering degree and management training a) What is the probability that an engineer (any engineer) will not have had any management 2, training? What is the probabilitv that an engincer with a mechanical engincering degree has had managcmcnt training? b) c) What is the probability an engineer will have some other degree or management training but not both? 8.50 x 11on in <Please answer with explanation this Statistic and Probability question ?
Answer #1
Management Training
Degree Yes No Total
Electrical 0.11 * 27 = 2.97 0.30 * 73 = 21.9 24.87
Mechanical 0.09 * 27 = 2.43 0.42 * 73 = 30.66 33.09
Software 0.02 * 27 = 0.54 0.21 * 73 = 15.33 15.87
Other 27 - (2.97 + 2.43 + 0.54) = 21.06 73 - (21.9 + 30.66 + 15.33) = 5.11 26.17
Total 0.27 * 100 = 27 100 - 27 = 73 100

(a) Probability = 73/100 = 0.73

(b) Probability = 2.43/100 = 0.0243

(c) Probability = 26.17/100 + 27/100 – 21.06/100 = 32.11/100 = 0.3211